# Trust No Bot: Discovering Personal Disclosures in Human-LLM Conversations in the Wild

### v1

@@ 1 Introduction
Our findings illuminate the many risks that are taken on by chatbot users. Whether these users are knowingly trading their privacy for chatbot access or are unaware that their data is being collected by chatbot companies (and the risks entailed by this collection), we believe these findings have strong implications for both chatbot designers and LLM researchers. We call for the design of appropriate nudging mechanisms to help users moderate their interactions Acquisti et al. (2017), as well as increased transparency from chatbot companies. We also call for further research in local, private models and increased attention from privacy and security scholars into these high-stakes conversations.
@@ 5 Discussion
The decision to self-disclose is contextual (Yang et al., 2019; Zhao et al., 2012; Li et al., 2018), and self-disclosure can be a sign of trust (Galegher et al., 1998) and growth in relationship intimacy (Altman & Taylor, 1973). When users self-disclose either about PII or about sensitive topics, this provides an indication of their level of trust with their interlocutor. However, based on the WildChat data, it is impossible to say whether each user perceives their interlocutor in this context as the chat tool, the underlying model, the parent company, the researchers who collected WildChat, or some combination of these. More research in human-computer interaction is needed to disentangle the users’ perceptions of their “relationships” with and trust in the chatbots.

### v2

@@ 1 Introduction
Our findings illuminate the many risks that are taken on by chatbot users. Whether these users are knowingly trading their privacy for chatbot access or are unaware that their data is being collected by chatbot companies (and the risks entailed by this collection), we believe these findings have strong implications for both chatbot designers and LLM researchers. We call for the design of nudging mechanisms to help users moderate their interactions Acquisti et al. (2017), as well as increased transparency from chatbot companies. Most importantly, we call for further research in local, private models and increased attention from privacy and security scholars into these high-stakes conversations.
@@ 5 Discussion
The decision to self-disclose is contextual (Yang et al., 2019; Zhao et al., 2012; Li et al., 2018), and self-disclosure can be a sign of trust (Galegher et al., 1998) and growth in relationship intimacy (Altman & Taylor, 1973). When users self-disclose either about PII or about sensitive topics, this provides an indication of their level of trust with their interlocutor, and evidence suggests that users may reciprocate “disclosures” made by dialog systems (Ravichander & Black, 2018). This kind of chatbot behavior can be explicitly designed to elicit users’ self-disclosures, which may be desirable for, e.g., supporting mental health or improving conversation quality (Lee et al., 2020; Ichino et al., 2022; Harmsen et al., 2023; Jo et al., 2024). Prior work has found that human-chatbot conversations can contain as much self-disclosure as human-human conversations, likely due to their perceived anonymity and lack of judgment compared to more trusted human interlocutors (Croes et al., 2024). Importantly, based on the WildChat data, it is impossible to say whether each user perceives their interlocutor in this context as the chat tool, the underlying model, the parent company, the researchers who collected WildChat, or some combination of these. More research in human-computer interaction is needed to disentangle users’ perceptions of their “relationships” with and trust in LLM-based chatbots like ChatGPT, and the design of chatbots should carefully balance features that encourage self-disclosure, application goals, and privacy concerns.
@@ 6 Related Work
Self-disclosure detection
Prior work on the detection of self-disclosures has focused on explicit disclosures statements (e.g., “My name is Maria,” “I live in Seattle”) (Bak et al., 2012; Ravichander & Black, 2018; Valizadeh et al., 2021; Reuel et al., 2022; Dou et al., 2023; Yang et al., 2024) rather than the implicit sensitive topics (e.g., discussion of sexually explicit topics without any personal statement) that we explore in this work. Methods for explicit self-disclosure detection have included topic modeling (Bak et al., 2014), LLM fine-tuning (Dou et al., 2023), multi-task models (Reuel et al., 2022), and LLM-based prompts (Yang et al., 2024). Much of this work has focused on the social dynamics of self-disclosure in online, public conversations on Twitter (Bak et al., 2012; 2014) and Reddit (Reuel et al., 2022; Dou et al., 2023; Valizadeh et al., 2021; Falenska et al., 2024; Yang et al., 2024), rather than the closed user-chatbot conversations that we study. Relevant exceptions include measurements of self-disclosure in therapy conversations (Shapira & Alfi-Yogev, 2024) and conversations with dialog systems and agents (Ravichander & Black, 2018; Cho et al., 2022); the latter studies revealed high rates of explicit self-disclosure, which our study (1) echoes in our detection of high rates of sensitive topics and (2) refines via task and topic categories.